OpenStudy (anonymous):

a cuboid container with a base length twice its width is to be made with 48 metres squared of metal. a) show that the height, h = 8/x - 2x/3, where x is the width of the base b) express the volume, V, in terms of x c) find the maximum volume

OpenStudy (campbell_st):

start by letting x = width so the length is 2x the areas top and bottom = 4x^2 (width = x, length = 2x) 2 faces front and back = 4xh (2x and height h) 2 faces left and right = 2xh (x and height h) 2 faces then using the information \[48 = 2x^2 + 4xh + 2xh~~~or~~~~ 48 = 2x^2 + 6xh\] then \[48 - 2x^2 = 6xh\] so \[h = \frac{48 - 4x^2}{6x} \] you can simplify it from here so the volume is \[V=lwh~~~or~~~ V = 2x \times x \times (\frac{8}{x} - \frac{2x}{3})\] you can simplify this... then the max volume find the 1st derivative and solve for x. This will be the value that gives the max volume hope it helps

OpenStudy (anonymous):

thank you very much for the reply. I understand it but I'm just confused as to how you got from \[48 - 2x^2 = 6xh\] to \[\frac{ 48 - 4x^2 }{6x}\] where did the 4x come from?

OpenStudy (campbell_st):

oops the 1st line should be for surface area 48 = 4x^2 + 6xh so 48 - 4x^2 = 6xh

OpenStudy (anonymous):

okay thanks alot! :) I hope this doesn't come off as stupid but is that the answer or do I need to use the quadratic formula to find x.. Im still kind of confused what the question is asking

OpenStudy (campbell_st):

well the volume becomes \[V = 2x^2 \times \frac{8}{x} - 2x^2 \times \frac{2x}{3}\] so the volume equation is \[x = 16x - \frac{2x^3}{3}\]

OpenStudy (campbell_st):

oops it starts with \[V = 2x \times x \times (\frac{8}{x} - \frac{2x}{3})\]